# Accelerating Mobile Language Model via Speculative Decoding and NPU-Coordinated Execution

### v3

@@ Abstract.
Enhancing on-device large language models (LLMs) with contextual information from local data enables personalized and task-aware generation, powering use cases such as intelligent assistants and UI agents. While recent developments in neural processors have substantially improved the efficiency of prefill on mobile devices, the token-by-token generation process still suffers from high latency and limited hardware utilization due to its inherently memory-bound characteristics. This work presents sd.npu, a mobile inference framework that integrates speculative decoding with dynamic hardware scheduling to accelerate context-aware text generation on mobile devices. The framework introduces three synergistic components: (1) adaptive execution scheduling, which dynamically balances compute graphs between prefill and decoding phases; (2) context-aligned drafting, which improves speculative efficiency through lightweight online calibration to current tasks; and (3) hardware-efficient draft extension, which reuses and expands intermediate sequences to improve processing parallelism and reduce verification cost. Experiments on multiple smartphones and representative workloads show consistent improvements of up to 3.8× in generation speed and 4.7× in energy efficiency compared with existing mobile inference solutions. Component-level analysis further validates the contribution of each optimization.
@@ 1. Introduction
Recent advances in large language models (LLMs) (Brown et al., 2020; OpenAI et al., 2024; Touvron et al., 2023; Guo et al., 2024), rising concerns over privacy, and rapid progress in hardware have fueled interest in on-device LLM. In particular, the integration of neural processing units (NPUs) (Qualcomm, 2024, 2025b) makes it feasible to run LLMs directly on mobile devices (Google, 2024; Huawei, 2024; Samsung, 2024; Apple, 2024a), enabling applications such as voice assistants and UI automation (Apple, 2024b; AI, 2024; Wen et al., 2024; Lee et al., 2024). To extend LLMs’ capabilities in handling complex tasks, context-augmented generation (CAG) has emerged as a key technique (Wang and Chau, 2024; Ouyang et al., 2025). As shown in Figure 1, CAG augments a user’s query with task-relevant contexts before inference to strengthen the knowledge grounding of LLM. With CAG, LLMs can, for example, summarize documents using user-uploaded files, generate responses by retrieving and integrating information from the chat history, and act as UI agents by reasoning over current UI screenshots captured from the operating system, significantly expanding their functionality.
Despite these benefits, on-device CAG still has high end-to-end latency, primarily from decoding (accounting for 73.5% of total inference, details in Section 2.1). When diving into the reasons, we find that the prefill bottleneck is largely eliminated by NPU-based chunked prefill (Xu et al., 2025b; Chen et al., 2025a; Xue et al., 2024) (only 2 s for 1–2K tokens context). In contrast, the decoding stage remains memory-bound, a scenario poorly suited for NPUs. Therefore, most existing work resorts to CPU execution, leading to high latency. Motivated by the NPU’s superior energy efficiency and lower contention, this paper raises the question whether NPUs can be effectively utilized to accelerate the decoding stage.
Figure 1. Workflow of context-augmented generation and comparison between sd.npu and related work :
To address this issue, we observe that CAG outputs exhibit strong similarity to both the interaction history and the augmented context (Section 2.2), making them well-suited for retrieval-based speculative decoding (R-SD), a technique originally developed for code generation, and in turn, to NPU acceleration. In many CAG tasks, generated content is semantically close to previously produced tokens and to the augmented context: for example, summarization and automated email reply often reuse spans from user-uploaded documents or earlier emails, while UI agents repeatedly emit fixed commands (e.g., “click set clock button”). These characteristics align naturally with R-SD, which proceeds in two stages: (1) drafting, where a retriever proposes candidate token sequences by selecting similar fragments from the augmented context or a local store (e.g., a code repository); and (2) verifying, where the LLM compares these drafts against its own next-token predictions and corrects discrepancies (see Section 2.2). By evaluating multiple drafts in parallel, R-SD shifts decoding from a memory-bound to a compute-bound workload, making it a strong fit for NPU acceleration.
In this paper, we present sd.npu, the first system to enable retrieval-based speculative decoding for efficient on-device CAG through end-to-end on-chip NPU offloading. During prefill, sd.npu employs chunked prefill-based NPU inference like existing approaches (Xu et al., 2025b; Chen et al., 2025a); during decoding, it dynamically switches to a compute graph optimized for decoding, performing retrieval-driven drafting and parallel verification. Designing sd.npu requires addressing the following three unique challenges of aligning retrieval-based speculative decoding with NPUs, which prior work has not explored:
$\bullet$ NPU execution: costly switching overhead for distinct prefill and decoding optimal compute graphs. Prefill favors large-shape graphs with long sequence inputs which degrade decoding performance, while decoding benefits from small-shape graphs that underperform for prefill. Since the NPU only supports inference on static shapes, achieving optimal end-to-end performance thus requires switching between prefill-optimal and decoding-optimal compute graphs, incurring a significant performance overhead. For instance, our preliminary evaluation on Qwen2.5-1.5B model shows this switching overhead incurs a 94.9% delay of end-to-end latency.
$\bullet$ Draft construction: task context and LLM outputs follow different lexical distributions, yielding a low acceptance ratio. In R-SD, the LLM only accepts tokens that exactly match its own prediction. Our preliminary evaluation shows that drafts constructed from the task context often have high semantic similarity to the target output yet exhibit notable lexical gaps (e.g., normalized Levenshtein distance >0.11>0.11 for Qwen2.5-1.5B). As a result, LLM often accepts only few tokens per draft, falling back to token-by-token decoding and underscoring the need for lexical distribution calibration.
$\bullet$ Draft verification: short or absent drafts fail to saturate NPU computation. Mobile NPUs excel at processing large batches or long sequences due to their weight-stationary design where short inputs obtain fewer benefits (Jouppi et al., 2017; Chen et al., 2025a). However, due to strict suffix-matching, existing R-SD methods often generate only a few candidate tokens or none at all. Such short sequences cannot fully exploit the NPU’s weight-stationary design, resulting in low verification throughput. For example, our experiments show that over 75% of the drafts fall below length 8, which is far below the verification-optimal length (64) and leads to 50% lower throughput.
To address these challenges, we propose three techniques:
$\bullet$ Progressive Graph Scheduling (Section 3.2). sd.npu deploys two specialized graphs and switches between them asynchronously to minimize overhead. The LLM is partitioned into blocks. During each round of chunked prefill, we switch to a decoding-optimized block graph while prefill computation continues in parallel. The switched decoding graph is invoked multiple times within the prefill stage to ensure result correctness. This design rests on two observations: (i) executing short-input graphs multiple times is equivalent to a single long-input execution, and (ii) graph loading can be overlapped with one-round chunked prefill computation. Consequently, sd.npu can transition to the decoding graph without additional load overhead, yielding a negligible graph-switching cost.
$\bullet$ In-Context Distribution Calibration (Section 3.3). Motivated by the observation that, during prefill, the model computes next-token logits for each context token that already align with its output distribution, sd.npu uses these logits to calibrate the task context by constructing a model-distributed token tree via depth-first search. In the subsequent retrieval-based drafting stage, sd.npu retrieves from this calibrated tree to mitigate lexical divergence between the context and the model’s expectations. Compared with costly precomputation or task-specific fine-tuning, this lightweight procedure incurs negligible overhead.
$\bullet$ NPU-Optimized Draft Reuse (Section 3.4). sd.npu extends draft construction by selectively reusing plausible tokens from previously rejected drafts based on a confidence-based strategy. This is motivated by the observation that rejected tokens may still capture correct semantics and be accepted in later decoding steps. By reusing such tokens, sd.npu increases effective draft length with useful content, improving NPU utilization and reducing verification costs.
@@
Our contributions are summarized as follows:
$\bullet$ We present the first study of inefficiencies in context-aware generation for mobile tasks, identifying key optimization opportunities of using retrieval-based speculative decoding.
$\bullet$ We design and implement sd.npu, an efficient NPU-optimized on-device context-augmented generation framework with three novel techniques, bridging the gap between retrieval-based speculative decoding and NPU offloading.
$\bullet$ We conduct comprehensive evaluations across four datasets, three smartphones, and three LLMs, demonstrating that sd.npu delivers consistent and significant performance gains in diverse mobile scenarios.
@@ 2. Background
2.1. On-device Context-Augmented Generation for Mobile Task
Context-augmented generation. With the rapid development of LLMs and increasing demands for privacy protection, there is a growing trend of deploying large models on mobile devices (Belcak et al., 2025; Navardi et al., 2025; Union, 2021; Kshetri, 2023; Eleftherakis et al., 2024). An emerging paradigm, context-augmented generation, leverages both the long-context capability of LLMs and the abundance of private on-device data, showing strong potential for supporting complex tasks such as personal assistant and UI automation (Wang and Chau, 2024; Ouyang et al., 2025; Park et al., 2025). Typically, context-augmented generation first exploits additional information, either directly provided by the user or retrieved from relevant sources, to enrich the prompt with task-specific contexts. Then, the normal LLM inference procedure is executed, including prefill and decoding stages. The enriched prompt can alleviate the limitations of the model’s built-in knowledge and enable it to handle a wider range of concrete tasks. For instance, in a summarization task, the user may supply a meeting transcript as context, and the LLM directly generates a concise summary from it. In another case, a personalized assistant retrieves a user’s past chat records to draft a reply in the user’s typical style, where the retrieved records provide the key context for generation.
LLMs on NPUs. As modern mobile SoCs increasingly integrate high-performance NPUs (Table 1), recent mobile LLM frameworks aim to offload LLM inference onto NPUs to improve efficiency. In practice, NPUs are mainly used for the prefill stage, which benefits from their high throughput on long input sequences, while decoding is still executed on CPUs or GPUs (Xu et al., 2025b; Chen et al., 2025a; Xue et al., 2024). Our experiments reveal that mobile CAG still suffers from high latency under this design, with decoding dominating end-to-end performance. For example, for Qwen2.5-1.5B (Qwen et al., 2025) running on a OnePlus 13, decoding accounts for 90.2% of latency in question answering (QA) tasks and 87.2% in summarization, even when prefill runs at over 1000 tokens/s (Xu et al., 2025b) (Figure 2). Simply offloading decoding to the NPU provides no advantage over CPU execution and often increases overall latency, as decoding is inherently memory-bound, processing one token at a time and leaving the NPU underutilized.
Figure 2. Comparison of prefill and decoding latency for several CAG tasks. Task details are provided in Section 4.1.
Table 1. Specifications of representative mobile NPUs.
@@ 2.1. On-device Context-Augmented Generation for Mobile Task
These findings are closely tied to the architectural characteristics of mobile NPUs. First, their systolic-array with weight-stationary design excels at long input sequences: weights are preloaded once and reused across tokens, greatly reducing memory traffic. This property aligns naturally with the prefill stage, which requires lots of computation over long prompts. Second, NPUs rely on static compute graphs with precompiled execution plans, where operator arrangements and tensor shapes are fixed in advance for efficiency. Naively reusing prefill-optimized graphs for decoding leads to severe inefficiencies. As a result, while NPUs eliminate prefill bottlenecks, they bring little benefit to decoding, which thus remains the dominant source of latency and energy consumption in on-device CAG.
2.2. Opportunity: Retrieval-based Speculative Decoding
A key observation is that CAG naturally inherits similarity from the request or the augmented context, making it well-suited for retrieval-based speculative decoding (Saxena, 2023; He et al., 2024; Hu et al., 2024; Somasundaram et al., 2024), an acceleration technique exploiting similarity between contexts and outputs. We find that mobile CAG tasks exhibit the same property, offering strong opportunities for acceleration via R-SD.
$\bullet$ History similarity in similar tasks: outputs from similar tasks resemble previous generations because their contexts are similar. For example, a UI automation agent often repeats commands like clicking the same button. Our evaluation on four UI automation task groups (Wen et al., 2024) with Qwen2.5-1.5B (Qwen et al., 2025) confirms this (Figure 3(a)): predictions within the same task show the highest semantic similarity (0.72), those within the same group moderate similarity (0.61), and across groups low similarity (0.38). Cosine similarity is computed using embeddings from all-MiniLM-L6-v2 (Wang et al., 2020).
$\bullet$ Context similarity at semantic level: Outputs align with the enriched context, which often includes direct clues for response (Wang and Chau, 2024; Ouyang et al., 2025; Park et al., 2025). For instance, assistants answering QA queries frequently extract information from documents. On a summarization dataset (Xia et al., 2024) with Qwen2.5-1.5B, we compute semantic similarity between context and predictions (via cosine similarity) and lexical overlap (via Levenshtein distance). Figure 3(b) shows that 90% of samples exhibit high semantic similarity (>0.76>0.76) but low lexical overlap (>0.11>0.11), reflecting a stylistic gap between context and outputs.
(a) History
(b) Context
Figure 3. Demonstration of similarities in context-augmented generation.
Figure 4. Workflow of retrieval-based speculative decoding.
R-SD exploits these similarities through two stages (Figure 4): (1) retrieval-based drafting, where relevant fragments are retrieved from history or context to draft candidate sequences. For example, tokens “need to” may retrieve drafts such as “click on” and “find the”; (2) verifying, where the LLM validates these candidates in parallel, accepting matches and discarding mismatches to guarantee correctness (Leviathan et al., 2023a). In the example where model expects “find the”, mismatched drafts like “click on” are discarded while the matched draft is accepted, extending the output to “find the Google”. The similarities in mobile CAG tasks offer promising draft candidates, enabling R-SD to efficiently generate multiple tokens in parallel. R-SD thus substantially raises the computation-to-memory ratio, shifting decoding from memory-bound to compute-bound where NPU excels.
Parallel to retrieval-based methods, model-based speculative decoding (M-SD) (e.g., EAGLE (Li et al., 2024c, b), Medusa (Cai et al., 2024)) trains auxiliary draft models to generate candidates, incurring extra training, storage, and inference overheads. R-SD avoids retraining by directly using contextual information of high similarity, achieving draft latencies under 10ms (Hu et al., 2024). Therefore, R-SD is more suitable for resource-constrained devices while M-SD suffers from higher latency, as shown in our evaluations (Section 4.2).
2.3. Challenges: Aligning Retrieval-based Speculative Decoding with NPU
Although R-SD offers a promising approach to accelerate CAG with NPU offloading, integrating it with NPU offloading poses several challenges.

### v4

@@ Abstract.
Performing Retrieval-Augmented Generation (RAG) directly on mobile devices is promising for data privacy and responsiveness but is hindered by the architectural constraints of mobile NPUs. Specifically, current hardware struggles with the variable workloads intrinsic to RAG: the transition between processing extensive contexts and generating tokens incurs significant overhead due to static graph constraints, while the memory-bound generation phase leaves computational resources underutilized.
[…]
@@ 1. Introduction
Retrieval-augmented generation (RAG) has become the de facto standard for grounding Large Language Model (LLM) outputs in user-specific data, enabling a new class of edge applications ranging from personalized document assistants to automated UI agents (usecase:ui_automation; work:mobile_rag1; work:mobile_rag2; work:mobile_rag3). As these applications increasingly handle sensitive personal data, RAG workloads are shifting towards on-device execution to ensure privacy and reduce cloud reliance (trend:slm1; trend:slm2; privacy1; privacy2; privacy3). In this setting, the retrieval phase is typically fast (e.g., <10<10 ms for local corpora (work:edge_rag; work:webanns)), leaving the computational heavy lifting to the LLM generation phase.
To sustain always-on RAG capabilities without draining the battery or stalling the user interface, the Neural Processing Unit (NPU) has emerged as the critical hardware accelerator. Unlike CPUs and GPUs that are burdened with OS services, application logic, and graphics rendering, NPUs are domain-specific architectures dedicated to tensor operations. They offer two distinct advantages: (i) superior energy efficiency, delivering significantly lower energy-per-token for Transformer workloads compared to general-purpose processors (work:mllm; work:heterollm; work:tman); and (ii) performance isolation, allowing background RAG inference to run concurrently with foreground tasks without contending for CPU/GPU resources (work:npu_sparse_attn). Driven by these benefits, recent systems have moved towards NPU-only inference (work:tman; work:ttc; work:npu_sparse_attn), executing the entire pipeline on the NPU to eliminate the high synchronization and data movement costs inherent in heterogeneous CPU-NPU co-scheduling (work:mllm; work:heterollm).
Despite these advantages, NPU-only on-device RAG inference still suffers from long decoding latency, driven by two critical challenges:
$\bullet$ Static graph architecture vs. dynamic RAG phases. Mobile NPUs rely on static compute graphs with fixed tensor shapes, buffer layouts, and execution plans. Yet RAG inference alternates between two computational extremes: prefill, which processes long, context-heavy prompts (suitable graph for 128/256 tokens input), and decoding, which autoregressively generates tokens one-by-one (suitable graph for 1 token input). The gap between these two regimes is too large that no single static graph can efficiently serve both. To fit within tight on-device memory budgets, existing systems keep only one graph resident at a time, necessitating an expensive teardown-and-load process when switching phases: unloading the large prefill graph and loading the decoding graph. We find that naive graph switching accounts for up to 29% of the decoding latency in on-device RAG settings.
$\bullet$ Matrix-oriented hardware vs. memory-bound decoding. While prefill saturates the NPU’s matrix units, decoding is memory-bound, leaving compute resources severely underutilized. Speculative Decoding (SD) (work:sd_is_lossless) is a standard remedy that enables parallel decoding through a draft-then-verify schema, but existing model-based SD methods (e.g., EAGLE (work:eagle)) require auxiliary models, leading to high mobile memory consumption and slow speculative inference on NPUs. For RAG specifically, retrieval-based SD (work:sam) is theoretically ideal: it drafts candidate tokens directly from context by first-token retrieval matching without requiring the extra cost of auxiliary draft models. However, we observe that off-the-shelf retrieval-based SD fails to saturate mobile NPUs due to two inefficiencies. To understand why, we introduce the concept of effective NPU utilization and decompose the NPU cycles spent during verification into three categories: (i) Valid: cycles spent verifying tokens that are eventually accepted; (ii) Wasted: cycles spent verifying tokens that are rejected; and (iii) Underutilized: idle capacity in the matrix engines when the workload is too small (low batch size/sequence length). Our analysis (§ 3) shows that existing approaches are dominated by the following inefficiencies:
  1. (i)
    Lexical Divergence causes Wasted Compute: retrieved context often matches the semantic intent but differs lexically from the LLM’s training distribution. This causes the verifier to reject most draft tokens. In our measurements, 22% of NPU cycles are wasted verifying drafts that are ultimately rejected.
  2. (ii)
    Short Drafts cause Underutilization: NPU matrix units (e.g., Qualcomm HMX (tool:qnn_sdk), Huawei Cube (tool:ascend)) rely on large operational intensity (large batch or sequence length) to hide memory latency and satisfy the minimal tile size (e.g., 32 tokens for HMX). However, retrieval-based drafts are often fragmented and short. We find that over 75% of drafts are shorter than 8 tokens, whereas mobile NPUs typically require sequences of 32+ tokens to approach peak utilization. This mismatch leaves more than 70% of the NPU’s compute capacity underutilized during verification steps.
Design We present sd.npu, the first system designed to enable efficient, end-to-end NPU offloading for on-device RAG. It bridges the gap between the NPU’s requirement for static, massive parallelism and the dynamic-shape, fragmented nature of RAG workloads by incorporating several techniques:
$\bullet$ Progressive graph scheduling (§ 4.2) addresses the monolithic loading stall caused by deferring graph switching to the final prefill stage. Instead of a sudden, blocking switch, we amortize the heavy I/O cost across the entire prefill phase by leveraging two key insights. (i) Decomposition: the monolithic LLM is partitioned into fine-grained blocks (subsets of Transformer layers), making the I/O payload of a single block small enough to be masked by several chunked prefill computation. (ii) Equivalence: we observe that repeated, iterative execution of a decoding-friendly block graph is mathematically equivalent to a single execution of a prefill-optimized graph. Based on these, sd.npu orchestrates a multi-iteration progressive transition. Specifically, by partitioning the LLM, we progressively orchestrates the block-by-block switching of the prefill graphs to their decoding-friendly counterparts; each block switching is meticulously designed to be perfectly overlapped with one chunked prefill operation. This ensures a seamless and zero-cost transition to the decoding phase. After the decoding graph for each block has been loaded, it is executed iteratively to process the corresponding chunk of the prefill input, thereby ensuring functional equivalence and maintaining the correct overall prefill output.
$\bullet$ NPU-tailored speculative decoding (§ 4.3, 4.4). Building upon retrieval-based speculative decoding, sd.npu transforms the ”wasted” and ”underutilized” NPU cycles into ”valid” generation throughput, incorporating two distinct techniques:
  1. (i)
    In-context distribution calibration (§ 4.3). To minimize wasted compute caused by lexical divergence, we must align the retrieved drafts with the model’s output distribution. Cloud-based solutions might perform offline fine-tuning on the knowledge base, but this is infeasible on edge devices due to battery, storage, and privacy constraints. Instead, sd.npu adopts a zero-cost, online calibration. We observe that during the prefill, the model inherently computes the logits for the provided context. sd.npu reuses these existing logits to construct a model-calibrated token tree. By drafting from this calibrated tree rather than the raw text, we reduce draft rejection without any additional computation or model training.
  2. (ii)
    NPU-optimized draft extension (§ 4.4). Minimizing underutilized compute caused by short drafts requires longer draft sequences that can saturate the NPU matrix units. Existing SD methods focus solely on the input side (drafting) and discard rejected tokens. However, we observe that rejection is often temporary: around 38.5% of tokens rejected in one step become valid in subsequent steps as the context evolves. Based on this insight, sd.npu implements a verification-aware draft extension. It recycles high-confidence rejected tokens from previous steps to extend the current draft. This not only recovers useful information but also artificially lengthens the verification batch, pushing the NPU into a higher-utilization regime.
@@
$\bullet$ Identify and characterize the key inefficiencies of end-to-end NPU offloading for RAG workloads, revealing optimization opportunities via speculative decoding.
$\bullet$ Design and implement sd.npu, a system that co-optimizes speculative decoding and NPU execution through two novel techniques, bridging the gap between compute-bound LLM decoding and NPU offloading.
$\bullet$ Conduct comprehensive evaluations across multiple devices, models, and tasks, demonstrating consistent latency and energy efficiency improvements in diverse edge environments.
@@ 2. Background
2.1. RAG for Mobile Applications
Retrieval‑augmented generation (RAG) has become an essential technique for on‑device intelligent applications (work:rag; work:mobile_rag1; work:mobile_rag2), such as mobile UI agents and personal assistants. By retrieving relevant external documents (e.g., UI information for automation and user archives for personalization) and injecting it into the prompt, RAG improves the accuracy and reliability of LLM‑driven tasks. Compared to cloud‑based inference, on‑device RAG offers strong privacy guarantees and consistent responsiveness, making it well suited for scenarios involving sensitive data or real‑time interactions (work:adaptive_rag). Recent efforts on retrieval acceleration have largely addressed its overhead, achieving latency of several microseconds for querying on GB-size corpus (work:webanns; work:edge_rag). Thus, our focus is the compute burden introduced by LLM inference, as the augmented context significantly lengthens prompts, amplifying the cost of both prefill and decoding (work:rag_amplify_cost1; work:rag_amplify_cost2). Fortunately, modern mobile SoCs are increasingly integrating neural processor units (NPUs) with high compute capacity, offering opportunities for hardware acceleration.
2.2. Mobile NPUs
Figure 1. RAG on NPU and NPU’s architecture.
Modern mobile NPUs are specialized accelerators designed to deliver high throughput and energy‑efficient execution for deep neural networks. Their compute cores are typically built around matrix computation units, which usually are implemented as systolic arrays with weight‑stationary designs (device:8gen3; device:8gen4). As illustrated in Figure 1, systolic arrays preload weight tiles into their processing elements and keep them stationary during computation. This weight‑stall paradigm amortizes the cost of weight movement and maximizes reuse across input activations, achieving high efficiency for large matrix multiplications. Meanwhile, the fixed systolic array constraints the minimal input dimensions to effectively saturate the processing units, as each tile in the array (e.g., 32×32 FP16 or INT8) corresponds to the fundamental execution unit of the NPU. To address this, the NPU runtime compiles operators into sequences of tile‑level matmul instructions, pads and partitions input tensors into tiles such that activations could be streamed through the systolic array with the optimal efficiency. Therefore, most mobile NPU SDKs require neural networks to be converted into static compute graphs with fixed tensor shapes, memory bindings, and operator layouts. Graph compilation can take several seconds and is generally performed offline (work:mllm). At runtime, the NPU executes these static graphs with minimal control‑flow support, which simplifies scheduling overhead by restricting dynamic shape changes and conditional execution.
2.3. Accelerating RAG on Mobile NPUs
@@
Table 1. Specifications of representative mobile NPUs.
As mobile SoCs increasingly integrate NPUs (Table 1), leveraging these accelerators for RAG has become a critical optimization direction. When deploying RAG systems on edge devices, existing inference engines typically adopt one of two execution strategies:
$\bullet$ NPU-CPU/GPU co-scheduling (e.g., mllm (work:mllm), HeteroLLM (work:heterollm), PowerInfer‑V2 (work:powerinfer2)): offloading only compute-intensive operations in prefill to NPU, while executing precision-intensive kernels and decoding on CPUs or GPUs.
$\bullet$ NPU-only execution (e.g., T-MAN (work:tman), test-time-compute NPU (work:ttc) and ExecuTorch (work:executorch)): offloading all computation to NPU through carefully designed quantization strategies.
Figure 2. Comparison on consumed energy per request across different execution backends (work:mllm; work:heterollm).
Among these, the NPU-only approach is particularly promising for sustainable RAG acceleration. As shown in Figure 2, NPU-only execution reduces energy consumption by eliminating frequent cross‑processor scheduling and by consolidating execution onto the most energy‑efficient compute unit (work:mllm; work:heterollm). Furthermore, it also minimizes resource contention with other general-purpose mobile workloads processed by CPU and GPU, such as user interaction handling and rendering (work:npu_sparse_attn). Therefore, this paper focuses on addressing the challenges within the NPU-only acceleration path.
However, realizing the full potential of NPU-only execution remains challenging. Prior work (e.g., test-time-compute NPU (work:ttc), T-MAN (work:tman)) primarily relies on aggressive quantization strategies (work:decdec), which necessitate invasive changes to NPU kernels and proprietary SDKs. Moving beyond kernel-centric approaches, we aim to optimize RAG performance through system-level innovations. Our optimizations address the scheduling and memory inefficiencies inherent in NPU runtimes and are designed to be orthogonal to specific quantization schemes, enabling additive performance gains atop these kernel-level baselines.
3. Challenges of RAG on NPU
Figure 1 depicts the NPU-based RAG workflow. Excluding the negligible retrieval overhead, execution comprises two distinct phases: compute-bound prefill and memory-bound decoding. During prefill, the variable-length context is partitioned into fixed-size chunks (e.g., 256 tokens) to accommodate the NPU’s static shape requirement (work:mllm; work:heterollm). However, this execution paradigm exposes two critical inefficiencies on resource-constrained mobile devices: (i) graph switching vs. padding dilemma (§ 3.1): due to strict memory constraints, mobile runtimes can typically reside only one active static graph. Given the drastic shape disparity between prefill (e.g., 256 tokens) and decoding (1 token), the system faces a prohibitive trade-off: either switch graphs, which incurs high latency from model reloading, or reuse the prefill graph via padding, which results in massive computation waste. (ii) NPU underutilization in decoding (§ 3.2): the decoding phase fails to saturate the NPU’s massive matrix parallelism. A typical widely-used technique to alleviate this by increasing concurrency is speculative decoding (SD) (work:sd_is_lossless; work:eagle2; work:rest). However, we find traditional SD still fails to saturate the NPU’s systolic array architecture due to a mismatch with the hardware cost model. Crucially, leaving NPU compute resources significantly idle compared to the dense prefill phase.
3.1. Overhead of Static Graph Management
[…]